OpenStudy (anonymous):

integrate mod (sin x-cosx ) 0 to p1/2

OpenStudy (anonymous):

\[\int_o^{\pi/4} -(sin x - cos x)dx + \int_{\pi/4}^{\pi/2} (\sin x - \cos x)dx \] Now you can do it.

OpenStudy (ash2326):

We have \[\int_0 ^{\pi/2} |\sin x - \cos x| dx\] We know that from 0 to \(\pi/4\) sin x < cos x so \[|\sin x - \cos x|= \cos x -\sin x\] and from \(\pi/4\) to\(\pi/2\) sin x> cos x so \[|\sin x - \cos x|= \sin x -\cos x\] We get \[\int_0 ^{\pi/2} |\sin x - \cos x| dx=\int_0 ^{\pi/4} -(\sin x - \cos x) dx+\int_{\pi/4} ^{\pi/2} (\sin x - \cos x) dx\] Can you do it now?

OpenStudy (anonymous):

can you tel me the anwser

OpenStudy (ash2326):

Just answer won't help you, Do you know how to integrate sin x and cos x?

OpenStudy (anonymous):

Consider this your lucky day... the answer is \(2(\sqrt 2 - 1)\).